OpenStudy (trent):

A car travels at an average speed of 56 miles per hour. How many miles does it travel in 3 hours and 15 minutes?

OpenStudy (anonymous):

Speed * time = distance

OpenStudy (anonymous):

3 hours 15 mintues= 3.25 hours

OpenStudy (anonymous):

15 minutes --------- = 1/4=.25 60 minutes

OpenStudy (mr.math):

\(56\times 3.25\) miles.

OpenStudy (trent):

thank you for you help